The speed of light in an unknown medium is measured to be 2.76x10^8 m/s 

  1. What is the index of refraction in this medium? 
  2. Does it match any of the material in the table? If not, is it similar to any object in the table? 
1
Expert's answer
2021-08-16T08:32:10-0400

The index of refraction

"n=\frac{c}{v}=\frac{3.00*10^8\:\rm m/s}{2.76*10^8\:\rm m/s}=1.09"
  1. No. It similar to air.